Wilson Lawn & Garden sells lawn and garden equipment and products, including 50-pound bags of fertilizer. Use the following information to determine the cost of goods sold and the value of the ending inventory of fertilizer bags under each of the designated cost flow assumptions. Wilson uses a periodic inventory system.

DATE

PER UNIT

TOTAL

June 1 Balance

300 bags at $10 per bag

$ 3,000

June 10 Sold

200 bags at $24 per bag

$ 4,800

June 11 Purchased

800 bags at $12 per bag

$ 9,600

June 15 Sold

500 bags at $25 per bag

$ 12,500

June 20 Purchased

500 bags at $13 per bag

$ 6,500

June 27 Sold

300 bags at $27 per bag

$ 8,100

LIFO   Cost of Goods Sold During June: ?   

June 30 Ending Inventory Value: ?   

FIFO   Cost of Goods Sold During June: ?

        June 30 Ending Inventory Value: ?

            Units                         LIFO $                        FIFO $

Beginning Inventory

+ Purchases                          ________                  _________                    _________

Available for Sale

-Ending Inventory                 ________                  _________                    _________

Cost of Goods Sold

                                                                LIFO $                        FIFO $

Sales

Less: Cost of Goods Sold               __________                _________

Gross Profit                                                            

Gross Profit %
Answer #1

Answer: Beginning Inventory + Purchases Available for Sale |- Ending Inventory Cost of Goods Sold Units 300 1,300 1,600 600 1

Calculation:

Units Purchased = 800 + 500 = 1,300 Bags
Cost of Goods Purchased = $9,600 + $6,500 = $16,100

Units available for Sale = Beginning Inventory Units + Units Purchased
Units available for Sale = 300 + 1,300 = 1,600 Bags

Units Sold = 200 + 500 + 300 = 1,000 Bags

Ending Inventory = Units available for Sale – Units Sold
Ending Inventory = 1,600 – 1,000 = 600 Bags

LIFO:
Cost of Ending Inventory = (300 * $10) + (300 * $12)
Cost of Ending Inventory = $6,600

Cost of Goods Sold = Cost of Goods available for Sale – Cost of Ending Inventory
Cost of Goods Sold = $19,100 - $6,600
Cost of Goods Sold = $12,500

FIFO:
Cost of Ending Inventory = (100 * $12) + (500 * $13)
Cost of Ending Inventory = $7,700

Cost of Goods Sold = Cost of Goods available for Sale – Cost of Ending Inventory
Cost of Goods Sold = $19,100 - $7,700
Cost of Goods Sold = $11,400